OpenStudy (anonymous):

Question 1.1. What is the solution of log4(2x 6) = 2? (Points : 1) 5 7 11 14

OpenStudy (ranga):

Something is missing between 2x and 6 within the parenthesis.

OpenStudy (anonymous):

It is subtraction between the 2x and 6...

OpenStudy (anonymous):

Non of your option are correct

OpenStudy (ranga):

\[If ~~\log_B(A) = C, ~then~~A = B^C\\If ~~\log_4(2x-6) = 2, ~then~~2x-6 = 4^2 = 16. ~~\text{ Solve for x.}\]

OpenStudy (anonymous):

Ohhhh, so it's 11! Thank you so so much!!

OpenStudy (anonymous):

I have another one I am confused about as well.. @ranga

OpenStudy (ranga):

\[log_B(A) - log_B(C) = log\left(\frac AC\right)\\log_{36}(x+4)-log_{36}(2x-14) = log_{36}\left(\frac{x+4}{2x-14}\right) = \frac 12\\\left(\frac{x+4}{2x-14}\right) = 36^{\frac 12} = \sqrt{36} =6.\quad \text{Solve for x.}\]